OpenStudy (anonymous):

A circle has its center at the origin, and (5, -12) is a point on the circle. How long is the radius of the circle? 5 12 13

OpenStudy (zzr0ck3r):

You want the distance from that point to the origin. Do you know how to find the distance between two points?

OpenStudy (anonymous):

 Simply use the triangle formula for perpendicular triangles, which says: \[c^2 = a^2 + b^2\] Where c is hyphotenus, a and b are sides. In here \[r^2 = 12^2 + 5^2 = 169\] \[r = \sqrt{169} = 13\]

OpenStudy (campbell_st):

well a simple method is to know the standard form of the equation of the circle \[(x - h)^2 + (y - k)^2 = r^2\] (h, k) is the centre... so substitute the given values then substitute the point this will allow you to calculate r^2 and then r... the radius... hope it helps

OpenStudy (anonymous):

thank you:)